The length of air column at which the first resonance was observed when a vibrating tuning fork was placed at the mouth of a resonance tube is 20 cm. Calculate
(i) The wavelength of the air column,

(ii) The frequency of the fork (assume the speed of sound in air is 340 m/s)

(iii) What is the frequency of the fork that would give resonance at 30 cm of the tube?
1
Expert's answer
2020-05-25T11:19:11-0400

i)


"\lambda=4l=4(0.2)=0.8\ m"

ii)


"f=\frac{v}{\lambda}=\frac{340}{0.8}=425\ Hz"

iii)


"f'=f\frac{l}{l'}=425\frac{20}{30}=283\ Hz"